OpenStudy (anonymous):

What is an easy way to divide 1/2 by (sq.rt.3)/2?

OpenStudy (saifoo.khan):

\[\frac12 \div \frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

lol caught you

OpenStudy (saifoo.khan):

\[\frac12 \times \frac{2}{\sqrt{3}}\]

OpenStudy (saifoo.khan):

\[\frac{1}{\cancel{2}} \times \frac{\cancel{2}}{\sqrt{3}}\]

OpenStudy (saifoo.khan):

\[\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

\[=3^{-\frac{1}{2}}\]

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

Satellite73, how did you come up with this?